OpenStudy (anonymous):

Determine the time necessary for P dollars to double when it is invested at interest rate r compounded monthly, daily. r =3.4%

OpenStudy (paxpolaris):

plug in 2P for A .034 for r solve for t

OpenStudy (anonymous):

then whats the difference between monthly and daily

OpenStudy (paxpolaris):

\[\Large 2\cancel P=\cancel P \left( 1 + {0.034\over12} \right)^{12t}\]and \[\Large 2\cancel P=\cancel P \left( 1 + {0.034\over365} \right)^{365 t}\]

OpenStudy (anonymous):

wait don't go yet let me solve it

OpenStudy (anonymous):

how do i get t out of an exponent

OpenStudy (paxpolaris):

take log of both sides

OpenStudy (anonymous):

i got 20.44 for monthly but it aid its wrong

OpenStudy (paxpolaris):

20.4155493 ... monthly 20.3876313 ... daily

OpenStudy (paxpolaris):

when you do 1+0.034/12, do NOT round

OpenStudy (anonymous):

ooooh! how do i know when to use a=pert and the one u gave me for this problem

OpenStudy (paxpolaris):

A=Pe^(rt) is when compounding continuously only

OpenStudy (anonymous):

ok thank you again how do u know all this?

OpenStudy (paxpolaris):

\[\Large A=P \left( 1+\frac rn \right)^{nt}\] is where you no compound n times in a year.

OpenStudy (paxpolaris):

*no = only

OpenStudy (anonymous):

thanks a lot